OpenStudy (anonymous):

Are the complex fractions x^2 - x- 20 / 4 / x-5/ 1- and 10/4/x-5/x^2 - x -20 equivalent???? how do i simplify each?

OpenStudy (anonymous):

there is something funny here. what is the 1- hanging out at the end?

OpenStudy (anonymous):

oh sorry! I mean 10

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

\[\frac{\frac{x^2-x-20}{4}}{\frac{x-5}{10}}\] like this one?

OpenStudy (anonymous):

no the x-1 is x-5

OpenStudy (anonymous):

oh sorry, i thought it said x-1

OpenStudy (anonymous):

and then divided by the other fraction

OpenStudy (anonymous):

ok so first we write \[\frac{x^2-x-20}{4}\times \frac{10}{x-5}\] then \[\frac{(x-5)(x+4)}{4}\times \frac{10}{x-5}\] then \[\frac{5(x+4)}{2}\]

OpenStudy (anonymous):

Okay :) and then what about teh next fraction?

OpenStudy (anonymous):

that one is \[\frac{10}{4}\times \frac{x^2-x-20}{x+5}\] so i guess they are identical.

OpenStudy (anonymous):

okay! wow, thaank you sooooo much satellite :D you're the best.

OpenStudy (anonymous):

Sorry, can you just explain why they are?